A 15-year $200000 mortagage has a fixed mortgage rate of 10 percent . Estimate the total mortagage payment and interest payment in the second month ?
Answer #1

Using financial calculator
Input:

N = 15*12 = 180

PV = -200000

I/Y = 10/12

Solve for PMT as 2149.21

Hence the monthly payment every month =

$2,149.21